OpenStudy (chillhill):

let h(x)= 6x/cos x. Find h'(0)

OpenStudy (studygurl14):

First find the derivative of h(x)

OpenStudy (studygurl14):

Then plug in 0 for x and solve

OpenStudy (chillhill):

how do you differentiate it?

zepdrix (zepdrix):

Hey ChillPill! Quotient rule, ya?\[\large\rm \left(\frac{6x}{\cos x}\right)'=\frac{(6x)'\cos x-6x(\cos x)'}{(\cos x)^2}\]

OpenStudy (chillhill):

won't it be -sinx for the differentiation of the denominator?

zepdrix (zepdrix):

\[\large\rm \left(\frac{6x}{\cos x}\right)'=\frac{(6x)'\cos x-6x(-\sin x)}{(\cos x)^2}\]Yes.